OpenStudy (anonymous):

What is 87 ÷ 10^4 ? A. 0.00087 B. 0.0087 C. 0.087 D. 0.87

OpenStudy (anonymous):

(87)/(10)(10)(10)(10)=87/1000=.087! Happy Studying!!!!!

OpenStudy (anonymous):

hence the answer must be c.

OpenStudy (anonymous):

87/(10*10*10*10) =?

OpenStudy (charlotte123):

10^4 ----> 10*10*10*10 = 10000 87/10000 = ?

OpenStudy (anonymous):

thnx everyone, and yes i have done it myself!!! but still!! thanks everyone!!! you all helped!!! :) :D

OpenStudy (charlotte123):

No problem! @ligiete